OpenStudy (anonymous):

find the exact values of sin 8pi/3 and cot -330 degrees

OpenStudy (anonymous):

{0.145742, 0.03023}

OpenStudy (anonymous):

Exact value infers a fraction, no calculator. The fraction 2/3 is exact compared to its decimal value of 0.66. 8 pi is just 4 revolutions of a circle...2 * PI * n so you can just use sin (pi/3) which is (3)^(1/2)/2 You could convert 330 degrees to radians by multiplying 1 rad/180 degrees. So you would get 33/18 radians.

OpenStudy (anonymous):

\[\sin (\frac{8\pi}{3}) = \frac{\sqrt{3}}{2}\] -330 degrees is the same thing as 360 + -330 which is 30 degrees.. Tangent of 30 degrees is \[\frac{\sqrt{3}}{3}\] Cot is \[\frac{1}{tanx}\] so \[\frac{1}{\frac{\sqrt{3}}{3}}\] or \[\frac{3\sqrt{3}}{3}\]